Please help...am so confused...taking stats online and have no teacher...thanks a bunch
The diameters of grapefruits in a certain orchard are normally distributed with a mean of 5.44 inches and a standard deviation of 0.41 inches. Show all work.
(A) What percentage of the grapefruits in this orchard is larger than 5.36 inches?
(B) A random sample of 100 grapefruits is gathered and the mean diameter is calculated. What is the probability that the sample mean is greater than 5.36 inches?

Answer by stanbon(57984)

The diameters of grapefruits in a certain orchard are normally distributed with a mean of 5.44 inches and a standard deviation of 0.41 inches. Show all work.
(A) What percentage of the grapefruits in this orchard is larger than 5.36 inches?
---
z(5.36) = (5.36-5.44)/0.41 = -0.1951
P(x > 5.36) = P(z > -0.1951) = 0.5774
===============================================
(B) A random sample of 100 grapefruits is gathered and the mean diameter is calculated. What is the probability that the sample mean is greater than 5.36 inches?
t(5.36) = (5.36-5.44)/[0.41/sqrt(100)] = -1.9512
---
P(x-bar > 5.36) = P(t > -1.9512 when df = 99) = 0.9731
================
Cheers,
Stan H.